please help me solve this.use matrix method to solve the system of linear equation.
x+2y+z=4
3x-4y-2z=2
5x+3y+5z=-1?

Answer by Alan3354(69443)

please help me solve this.use matrix method to solve the system of linear equation.
x+2y+z=4
3x-4y-2z=2
5x+3y+5z=-1?
---------------
If you mean determinants, just look it up on google, or some other search engine, or wikipedia, or youtube.